OpenStudy (anonymous):

What is the solution to the equation 7 3x ≈ 9

OpenStudy (anonymous):

is that 73x = 9 ? x=9/37

OpenStudy (anonymous):

nope

OpenStudy (zepp):

7*3x=9?

OpenStudy (zepp):

3*7=21 21x=9 Can you take it from there?

OpenStudy (anonymous):

ok i solved 73x=9 for u cant you solve this one on your own

OpenStudy (anonymous):

try harder man this is so easy and you know it